# and Regional Knowledge of Kazakhstan

### v1

@@ 4.2 Results and Analysis
Mistral-7B-Instruct-v0.3 36.5 44.7 41.8 30.0 40.6 38.2
Mistral-7B-v0.3 31.5 37.1 37.9 27.3 34.6 33.3
@@
bloomz-1b7 24.2 22.7 22.7 23.0 25.8 23.4
bloomz-3b 24.2 23.5 23.5 22.1 25.2 23.6
bloomz-7b 24.0 24.0 24.2 23.9 22.5 24.0
@@
Gemma-2-27b-IT 57.0 60.3 54.6 39.1 48.3 54.0
Gemma-2-9b 50.8 57.4 53.8 36.2 42.6 50.2
Gemma-2-9b-IT 50.6 52.8 49.1 35.8 44.3 48.1
@@
Llama3.1-70b-instruct 50.6 47.5 47.9 33.9 45.6 46.5
Llama3.1-8b 36.6 43.8 42.4 28.9 38.3 38.0
Llama3.1-8b-instruct 41.6 47.4 45.4 30.4 39.3 41.6
@@
DeepSeek V3 78.7 82.3 77.0 61.2 65.1 75.6
YandexGPT 53.7 69.8 52.2 42.6 57.0 54.9
Table 4: Performance of different models on the Kazakh MMLU benchmark across different subject categories using English prompt.
@@
Table  4 presents the average accuracy for each subject area across 26 models on the KazMMLU using English prompts. The performance analysis reveals several notable patterns. GPT-4o and DeepSeek V3 emerge as the top performers, achieving remarkably similar average scores of 75.7% and 75.6% respectively, outperforming other models. Among the open-source models, the Llama3.1 and Gemma families demonstrated strong performance. Llama3.1-70b achieved the highest average score (55.2%), followed by the Gemma-2-27b instruction-tuned model (54.0%) and its base variant (52.7%). Interestingly, the impact of instruction tuning varies across model families - while Gemma-2-27b-it showed a slight improvement over its base model (+1.3%), Llama3.1-70b-instruct performed worse than its base variant (-8.7%). Consistently across all models, the Language category seems to be the most challenging, with scores lower than other categories.
@@
In Figure 6, we compared model performance across two languages: Kazakh and Russian. To ensure a fair comparison, we only considered the High School level subjects and excluded the Professional & University level tasks because they are not available in Kazakh. The results indicate that GPT-4o achieves the highest accuracy in Kazakh, scoring 76.90%, while DeekSeek performs best in Russian with an accuracy of 81.8%. Llama 3.1-70B and Gemma 2-27B show lower but comparable results, with a slight advantage in Russian over Kazakh. Overall, models tend to perform slightly better in Russian than in Kazakh, which could be due to differences in training data availability, language complexity, or tokenization differences.
@@
Results in Figure 7 indicate that GPT-4o performs approximately the same across both education levels. Similarly, DeepSeek V3 maintains a balanced performance, but with a slight preference towards High School. In contrast, the open-source models Llama 3.1-70B and Gemma 2-27B show considerably lower accuracy and a pronounced gap between education levels. Llama 3.1-70B achieves 57.7% in High School but drops to 53.2% in Professional & University. These results suggest that while proprietary models generalize well across different subject complexities, open-source models struggle more with specialized university-level knowledge.
@@
Table 5 presents the accuracy of LLaMA3-70B, Gemma-2-27B-IT, and DeepSeek V3 on negation-sensitive subjects, comparing performance with and without negation. The results indicate that DeepSeek V3 consistently outperforms both LLaMA3-70B and Gemma-2-27B-IT, demonstrating greater resilience to negation-based reasoning challenges. To systematically analyze the impact of negation, we employed a negation phrase filtering method inspired by ArabicMMLU Koto et al. (2024). Specifically, we identified questions containing common negation phrases in Kazakh: жо (no), емес (is not), болмайды (not allowed), жарамайды (prohibited), невозможно (impossible), не (not), нельзя (forbidden). . After applying this filtering, we obtained a total of 2,554 negation-related questions. To validate our filtering accuracy, we randomly sampled 100 questions and manually inspected them. The detection accuracy exceeded 92%, confirming the reliability of our filtering process. For LLaMA3-70B and Gemma-2-27B-IT, accuracy generally decreases on questions containing negation, suggesting a negative impact on reasoning capabilities. Notably, LLaMA3-70B exhibits a larger drop in accuracy, particularly in Reading Literacy, where performance declines from 57.1% to 50.0%. Meanwhile, Gemma-2-27B-IT demonstrates greater robustness in certain cases, showing less fluctuation in performance across negation and non-negation settings.
Figure 6: LLM Performance across different languages(only at the high-school level).
Figure 7: LLM Performance across different education levels
@@
Jurisprudence (University) Llama3-70B 56.2 55.2
@@
Law (High School) Llama3-70B 60.8 59.0
@@
Reading Literacy (High School) Llama3-70B 57.1 50.0
@@
Philosophy and Psychology (University) Llama3-70B 55.9 56.6
@@
Model Confidence
We analyze whether the evaluated models, including Gemma-2.9B-IT, Qwen-2.5-7B-Instruct, and Llama3.1-70B, are well-calibrated when answering KazMMLU questions by comparing the probability of the correct answers with the actual accuracy for each subject and level combination. The answer probability is obtained through softmax normalization across the available candidate answers. In Figure 8, we observe that the evaluated models exhibit a strong correlation between confidence and accuracy, with correlation scores exceeding 0.9. This indicates that models with higher confidence generally produce more accurate predictions.
@@
Additionally, we investigate the correlation between model confidence and question length in Figure 12. The results show that question length has minimal impact on model confidence, as evidenced by the weak correlation scores across all evaluated models. For Qwen-2.5-7B-Instruct ($r=0.29$) and Llama3.1-70B ($r=0.28$) exhibit a mild positive correlation, suggesting that longer questions slightly increase model confidence. However, the effect remains weak overall, implying that confidence calibration remains relatively stable across different question lengths.
@@ 5 Conclusion and Future Work
We introduced KazMMLU, the first large-scale multi-task language understanding dataset designed to evaluate real-world knowledge in Kazakhstan’s bilingual setting. Through experiments with over 23K multiple-choice questions spanning various subjects and education levels, we observed that models perform much better in Russian than in Kazakh, with proprietary models such as GPT-4o and DeepSeek V3 achieving the highest accuracy. KazMMLU provides a bilingual (Kazakh and Russian) evaluation framework tailored to Kazakhstan’s educational and professional landscape, distinguishing itself from previous multilingual benchmarks. Unlike SIGTURK, which focused on Kazakh NLP tasks, and INCLUDE, which lacks Kazakhstan-specific content, KazMMLU enables a localized and comprehensive assessment of LLMs in Kazakhstan. Future research directions include extending KazMMLU to multimodal evaluations, improving reasoning-based question assessments, and mitigating biases in data sources. We hope that this benchmark will encourage further development of high-performance LLMs for Kazakh and other low-resource languages.
@@
While we are confident that our benchmark will significantly advance the development of Kazakh LLMs, it is important to acknowledge certain limitations that need to be addressed in future research. While we are confident that our benchmark will significantly advance the development of Kazakh LLMs, it is important to acknowledge certain limitations that need to be addressed in future research. We outline these limitations as follows:
@@
KazMMLU primarily evaluates static model performance on pre-defined questions, which may not fully capture how models generalize to dynamic, real-world language use. Exploring benchmarks that assess interactive and adaptive reasoning, as well as domain adaptation, could enhance our understanding of model capabilities in evolving contexts. By addressing these limitations, future research can further refine the evaluation of Kazakh LLMs, ensuring more robust, fair, and practically useful language models for Kazakhstan and beyond.
@@
We adhered to the internal policies of web resources while scraping data and included only publicly available information verified by authorities. All human subjects in our study provided informed consent, were fully aware of the study’s objectives, and had the right to withdraw at any time. They were also appropriately compensated as part of their job.
@@ B.1 Additional Multiple-Choice Question Example
Figure 12: Correlation between model confidence and question length
@@ B.2 Breakdown of Dataset Sources
B.3 Additional Evaluation Details
For further insights into our evaluation framework, we provide additional examples and implementation details. Table 4 summarizes our experimental settings.
@@ B.3 Additional Evaluation Details
Mistral-7B-Instruct-v0.3 33.3 41.2 37.6 28.0 31.9 34.9
Mistral-7B-v0.3 28.7 29.2 31.2 24.6 25.2 28.6
@@
bloomz-1b7 23.3 21.1 21.1 23.3 20.1 22.3
bloomz-3b 23.7 21.4 20.9 21.9 17.8 22.3
bloomz-7b 22.7 21.0 22.5 21.1 21.5 22.0
@@
gemma-2-27b-it 58.7 63.1 57.1 40.9 46.3 56.1
gemma-2-9b 47.3 54.6 52.4 33.6 39.3 47.4
gemma-2-9b-it 50.5 51.9 48.3 35.2 41.9 47.6
@@
Llama3.1-70b-instruct 56.3 57.4 55.5 39.5 50.3 53.4
Llama3.1-8b 33.7 38.4 39.0 28.0 31.5 34.7
Llama3.1-8b-instruct 36.9 39.5 39.9 28.0 31.5 36.4
@@
DeepSeek V3 78.5 81.8 77.1 61.4 65.3 72.8
YandexGPT 39.3 50.7 48.1 44.0 34.6 44.4
Table 7: Performance of different models on the Kazakh MMLU benchmark across different subject categories using Kazakh prompt.

### v2

@@ 4.2 Results and Analysis
Mistral-7B-Instruct-v0.3 37.6 47.4 43.3 30.0 40.6 41.0
Mistral-7B-v0.3 32.6 38.1 36.5 27.4 34.6 34.6
@@
Bloomz-1.7B 24.0 22.4 23.5 23.0 25.8 23.4
Bloomz-3B 24.4 23.9 22.8 22.1 25.2 23.7
Bloomz-7B 23.8 24.1 23.4 23.9 22.5 23.8
@@
Gemma-2-27B-IT 57.3 60.5 63.2 39.1 48.3 57.4
Gemma-2-9B 51.4 57.8 55.5 36.2 42.6 52.3
Gemma-2-9B-IT 50.8 53.0 48.3 35.8 44.3 49.1
@@
Llama3.1-70B-instruct 51.2 48.1 51.3 33.9 45.6 48.3
Llama3.1-8B 37.9 44.6 41.7 28.9 38.3 39.7
Llama3.1-8B-instruct 43.0 47.8 50.6 30.4 39.3 44.6
@@
[…]
DeepSeek V3 77.6 81.3 78.9 61.2 65.1 76.9
YandexGPT 54.8 70.6 63.7 42.6 57.0 60.2
Table 4: Performance of different models on the Kazakh MMLU benchmark across different subject categories using English prompt. “Average” means the average across all subject areas in KazakhMMLU.
@@
Table  4 presents the average accuracy for each subject area across 27 models on the KazMMLU using English prompts. The performance analysis reveals several notable patterns. GPT-4o and DeepSeek V3 emerge as the top performers, achieving remarkably similar average scores of 76.6% and 76.9% respectively, outperforming other models. Gemma-2-27B instruction-tuned model achieved the highest open-source score (57.4%), followed by Llama3.1-70B (56.2%) and its base variant (55.7%). Interestingly, the impact of instruction tuning varies across model families - while Gemma-2-27B-IT showed a slight improvement over its base model (+1.7%), Llama3.1-70B-instruct performed worse than its base variant (-7.9%). However, for smaller models, instruction tuning appears more beneficial, as seen in Llama3.1-8B-instruct outperforming its base model by 4.9% (44.6% vs 39.7%).
Consistently across all models, the Language category seems to be the most challenging, with scores lower than other categories.
@@
In Figure 6, we compared model performance across two languages: Kazakh and Russian. To ensure a fair comparison, we only considered the High School level subjects and excluded the Professional & University level tasks because they are not available in Kazakh. The results indicate that GPT-4o achieves the highest accuracy in Kazakh, scoring 76.90%, while DeekSeek performs best in Russian with an accuracy of 81.4%. Llama3.1-70B and Gemma-2-27B-IT show lower but comparable results, with a slight advantage in Russian over Kazakh. Overall, models tend to perform slightly better in Russian than in Kazakh, which could be due to differences in training data availability, language complexity, or tokenization differences.
@@
Results in Figure 7 indicate that GPT-4o performs approximately the same across both education levels. Similarly, DeepSeek V3 maintains a balanced performance, but with a slight preference towards High School. In contrast, the open-source models Llama3.1-70B and Gemma-2-27B-IT show considerably lower accuracy and a pronounced gap between education levels. Llama3.1-70B achieves 57.7% in High School but drops to 53.2% in Professional & University. These results suggest that while proprietary models generalize well across different subject complexities, open-source models struggle more with specialized university-level knowledge.
Figure 6: LLM Performance across different languages (only at the high school level).
@@
Table 5 presents the accuracy of Llama3.1-70B, Gemma-2-27B-IT, and DeepSeek V3 on negation-sensitive subjects, comparing performance with and without negation. The results indicate that DeepSeek V3 consistently outperforms both Llama3.1-70B and Gemma-2-27B-IT, demonstrating greater resilience to negation-based reasoning challenges.
@@
Jurisprudence (University) Llama3.1-70B 56.2 55.2
@@
Law (High School) Llama3.1-70B 60.8 59.0
@@
Reading Literacy (High School) Llama3.1-70B 57.1 50.0
@@
Philosophy and Psychology (University) Llama3.1-70B 55.9 56.6
@@
Figure 7: LLM Performance across different education levels.
@@
To systematically analyze the impact of negation, we employed a negation phrase filtering method inspired by ArabicMMLU Koto et al. (2024). Specifically, we identified questions containing common negation phrases in Kazakh: жоқ (no), емес (is not), болмайды (not allowed), жарамайды (prohibited), невозможно (impossible), не (not), нельзя (forbidden) . After applying this filtering, we obtained a total of 2,554 negation-related questions. To validate our filtering accuracy, we randomly sampled 100 questions and manually inspected them. The detection accuracy exceeded 92%, confirming the reliability of our filtering process.
For Llama3.1-70B and Gemma-2-27B-IT, accuracy generally decreases on questions containing negation, suggesting a negative impact on reasoning capabilities. Notably, Llama3.1-70B exhibits a larger drop in accuracy, particularly in Reading Literacy, where performance declines from 57.1% to 50.0%. Meanwhile, Llama3.1-70B demonstrates greater robustness in certain cases, showing less fluctuation in performance across negation and non-negation settings.
Model Confidence
We analyze whether the evaluated models, including Gemma-2.9B-IT, Qwen-2.5-7B-Instruct, and Llama3.1-70B, are well-calibrated when answering KazMMLU questions by comparing the probability of the correct answers with the actual accuracy for each subject and level combination. The answer probability is obtained through softmax normalization across the available candidate answers. In Figure 8, we observe a strong correlation between model confidence and accuracy, with correlation scores exceeding 0.9 across the evaluated models. This indicates that models with higher confidence generally produce more accurate predictions.
Additionally, we investigate the correlation between model confidence and question length in Figure 12. The results show that question length has minimal impact on model confidence, as evidenced by the weak correlation scores across all evaluated models. Qwen-2.5-7B-Instruct ($r=0.29$) and Llama3.1-70B ($r=0.28$) show a mild positive correlation, suggesting that longer questions slightly increase model confidence. However, the effect is weak overall, indicating that confidence calibration remains relatively stable across different question lengths.
@@ 5 Conclusion and Future Work
We introduced KazMMLU, the first large-scale multi-task language understanding dataset designed to evaluate real-world knowledge in Kazakhstan’s bilingual setting. Through experiments with over 23K multiple-choice questions spanning various subjects and education levels, we observed that models perform much better in Russian than in Kazakh, with proprietary models such as GPT-4o and DeepSeek V3 achieving the highest accuracy.
KazMMLU provides a bilingual (Kazakh and Russian) evaluation framework tailored to Kazakhstan’s educational and professional landscape, distinguishing itself from previous multilingual benchmarks. Unlike SIGTURK, which focused on Kazakh NLP tasks, and INCLUDE, which lacks Kazakhstan-specific content, KazMMLU enables a localized and comprehensive assessment of LLMs in Kazakhstan.
Future research directions include extending KazMMLU to multimodal evaluations, improving reasoning-based question assessments, and mitigating biases in data sources. We hope that this benchmark will encourage further development of high-performance LLMs for Kazakh and other low-resource languages.
@@
While we are confident that our benchmark will significantly advance the development of Kazakh LLMs, it is important to acknowledge certain limitations that need to be addressed in future research. We outline these limitations as follows:
@@
KazMMLU primarily evaluates static model performance on pre-defined questions, which may not fully capture how models generalize to dynamic, real-world language use. Exploring benchmarks that assess interactive and adaptive reasoning, as well as domain adaptation, could enhance our understanding of model capabilities in evolving contexts.
By addressing these limitations, future research can further refine the evaluation of Kazakh LLMs, ensuring more robust, fair, and practically useful language models for Kazakhstan and beyond.
@@
We adhered to the internal policies of web resources while scraping data and included only publicly available information verified by authorities.
All human subjects in our study provided informed consent, were fully aware of the study’s objectives, and had the right to withdraw at any time. They were also appropriately compensated as part of their job.
@@ B.1 Additional Multiple-Choice Question Example
Figure 12: Correlation between model confidence and question length.
@@ B.2 Breakdown of Dataset Sources
B.3 Evaluation Results using Kazakh Prompt
Table 7 presents the experimental results using Kazakh prompts across various models.
@@ B.3 Evaluation Results using Kazakh Prompt
Mistral-7B-Instruct-v0.3 34.9 43.3 38.3 27.9 31.9 37.3
Mistral-7B-v0.3 29.5 29.3 27.7 24.6 25.2 28.4
@@
Bloomz-1.7B 22.8 20.6 23.1 23.3 20.1 22.2
Bloomz-3B 23.2 20.6 23.6 21.9 17.8 22.2
Bloomz-7B 22.4 20.0 21.3 21.1 21.5 21.3
@@
Gemma-2-27B-it 59.3 62.6 63.3 40.9 46.3 58.8
Gemma-2-9B 47.9 54.0 54.0 33.7 39.3 49.3
Gemma-2-9B-it 50.4 51.6 48.8 35.2 41.9 48.5
@@
Llama3.1-70B-instruct 56.7 56.2 59.8 39.5 50.3 55.2
Llama3.1-8B 34.3 39.5 38.6 28.0 31.5 36.0
Llama3.1-8B-instruct 37.7 40.2 42.6 28.0 31.5 38.3
@@
qwen-2.5-7B-instruct 48.0 52.3 56.6 33.2 43.3 49.4
[…]
DeepSeek V3 45.4 48.0 47.8 37.4 33.3 45.5
Table 7: Performance of different models on the Kazakh MMLU benchmark across different subject categories using Kazakh prompt. "Average" means the average across all subject areas in KazakhMMLU.